# XiHe: A Data-Driven Model for Global Ocean Eddy-Resolving Forecasting

### v2

@@ 5.4 Ocean Mesoscale Eddy Evaluation with Satellite Observation
As we all know, forecasting eddy trajectory is a big challenge. In order to further test XiHe’s capability in forecasting the mesoscale eddy trajectory, we analyze the number of overlapping eddy trajectories of XiHe’s 7-day forecasting data and GLORYS12’s reanalysis data with CMEMS observations. We select the eddy trajectories with lifetimes larger than 60 days that appeared in both CMEMS data and GLORYS12 reanalysis data from Jan. 1, 2019, to Dec. 31, 2020. We only find 6 overlapping trajectories as shown in Fig. 14. Following the same methodology, we find 11 pairs of overlapping trajectories that appeared in both CMEMS data and XiHe’s forecasting data. It shows the good ability of XiHe in forecasting the mesoscale eddy trajectory. XiHe can promise a stronger eddy forecasting capacity with the quality improvement of the reanalysis data in the future.
@@ 6 Conclusion and Discussion
In this paper, we propose the first data-driven $1/12^{\circ}$ resolution global ocean eddy-resolving forecasting model XiHe, which is trained on 25-year ocean reanalysis data. An ocean-land masking mechanism is proposed to make XiHe focus more on the global ocean information by excluding the impact of land on the model training. Furthermore, a novel ocean-specific block containing both local and global SIE modules is designed to capture the inherent oceanic spatial information. Extensive experiments are conducted under satellite observations, in situ observations, and the IV-TT Class 4 evaluation framework from January 2019 to December 2020. The experimental results show that XiHe achieves stronger forecast performance in all testing variables than existing leading operational numerical GOFSs including PSY4, GIOPS, BLK, and FOAM with thousands of times speeding up. Notably, the 60-day ocean current forecasting of XiHe is even better than that of PSY4’s 10-day forecasting. The ACOR values of temperature and salinity profiles of XiHe’s 20-day forecast are 0.6177 and 0.5260, which are better than both PSY4’s and GIOPS’s 10-day forecasts. In addition, XiHe is able to forecast the large-scale circulation and the mesoscale eddies.
@@ Acknowledgments
This research is partially supported by National Natural Science Foundation of China (Grant No. 62372460, 42276205, 42030405), National Key R&D Program of China (Grant No. 2021YFC3101500), Hunan Provincial Natural Science Foundation of China (2023JJ10053), and the science and technology innovation Program of Hunan Province (2022RC3070). The authors would like to thank the Intercomparison and Validation Task Team of the OceanPredict Community for providing the observations and the forecasts of the oprational oceanography forecasting systems.

### v4

@@ 5.4 Ocean Mesoscale Eddy Evaluation with Satellite Observation
As we all know, forecasting eddy trajectory is a big challenge. In order to further test XiHe’s capability in forecasting the mesoscale eddy trajectory, we analyze the number of overlapping eddy trajectories of XiHe’s forecasting data and GLORYS12’s reanalysis data with CMEMS observations. We select the eddy trajectories with lifetimes larger than 60 days that appeared in both CMEMS data and GLORYS12 reanalysis data from Jan. 1, 2019, to Dec. 31, 2020. We only find 6 overlapping trajectories as shown in Fig. 14. Following the same methodology, we find 11 pairs of overlapping trajectories that appeared in both CMEMS data and XiHe’s forecasting data.
It shows the good ability of XiHe in forecasting the mesoscale eddy trajectory. XiHe can promise a stronger eddy forecasting capacity with the quality improvement of the reanalysis data in the future.
@@ 6 Conclusion and Discussion
In this paper, we propose the first data-driven $1/12^{\circ}$ resolution global ocean eddy-resolving forecasting model XiHe, which is trained on 25-year ocean reanalysis data. An ocean-land masking mechanism is proposed to make XiHe focus more on the global ocean information by excluding the impact of land on the model training. Furthermore, a novel ocean-specific block containing both local and global SIE modules is designed to capture the inherent oceanic spatial information. XiHe achieves stronger forecast performance in all testing variables than existing leading operational numerical GOFSs including PSY4, GIOPS, BLUElink OceanMAPS, and FOAM with thousands of times speeding up. Notably, the 60-day ocean current forecasting of XiHe is even better than that of PSY4’s 10-day forecasting. The ACC values of temperature and salinity profiles of XiHe’s 20-day forecast are 0.6177 and 0.5260, which are better than both PSY4’s and GIOPS’s 10-day forecasts. In addition, XiHe is able to forecast the multi-scale ocean processes, including the large-scale circulation and the mesoscale eddies.
@@ Acknowledgments
X.W., H.Z.W., G.H.W, and P.Q.W. were supported by National Natural Science Foundation of China (Grant No. 62372460, 42276205, 42288101, 42030405, 42306040). W.M.Z. was supported by National Key R&D Program of China (Grant No. 2021YFC3101500). H.Z.W. and X.W. were supported by Hunan Provincial Natural Science Foundation of China (2023JJ10053, 2024JJ4042). K.J.R. were supported by the Science and Technology Innovation Program of Hunan Province (2022RC3070). The authors would like to thank the Intercomparison and Validation Task Team of the OceanPredict Community for providing the observations and the forecasts of the operational oceanography forecasting systems.